OpenStudy (anonymous):

Need help solving ln(x+6)-lnx=ln79

OpenStudy (anonymous):

would that be ln(x+6)/lnx=ln79?

OpenStudy (mertsj):

\[\ln a - \ln b = \ln \frac{a}{b}\]

OpenStudy (anonymous):

what would i do after that?

OpenStudy (unklerhaukus):

once you have ln[something] = ln[something else] raise both sides to the base of the log e^ln[something] = e^ln[something else] then the logs and the base will cancle something = something else

OpenStudy (anonymous):

I'm sorry but that really confused me :o

OpenStudy (anonymous):

so i would do lne^(x+6)/(x)=lne79??

OpenStudy (mertsj):